# ValuesRAG: Enhancing Cultural Alignment Through Retrieval-Augmented Contextual Learning

### v2

@@ Abstract
Cultural values alignment in Large Language Models (LLMs) is a critical challenge due to their tendency to embed Western-centric biases from training data, leading to misrepresentations and fairness issues in cross-cultural contexts. Recent approaches, such as role-assignment and few-shot learning, often struggle with reliable cultural alignment as they heavily rely on pre-trained knowledge, lack scalability, and fail to capture nuanced cultural values effectively. To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with In-Context Learning (ICL) to integrate cultural and demographic knowledge dynamically during text generation. Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual. Subsequently, we curate several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries. ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided. Notably, ValuesRAG demonstrates an accuracy of 21% improvement over other baseline methods, highlighting its potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
1 Introduction
The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  Singh et al. (2024); Kharchenko et al. (2024); Hu et al. (2024). Predominantly trained on Western data sources Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts. These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024); Huang et al. (2024). Despite ongoing efforts to address these issues, existing strategies often fall short. While some countries have developed localized LLMs, such as China’s ERNIE Sun et al. (2021), ChatGLM GLM et al. (2024), DeepSeek Liu et al. (2024a), and South Korea’s HyperCLOVA Yoo et al. (2024), these models also exhibit biases inherited from their respective training datasets. As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024).
@@ 1 Introduction
Recent studies have proposed several approaches, such as role-assignment approaches Tao et al. (2024) and few-shot learning techniques  Choenni and Shutova (2024), to mitigate these cultural biases. However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
@@
To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL) to dynamically incorporate cultural knowledge during text generation (see Figure 1). Our framework leverages the World Values Survey (WVS) dataset Haerpfer et al. (2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies. Specifically, we first generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel. After constructing the knowledge base, we retrieve the top $100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected. Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies reasoning grounded in the retrieved values to generate the final answer to the question.
@@
We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method Tao et al. (2024), (3) few-shot learning Choenni and Shutova (2024), and (4) a hybrid method combining (1) and (2). To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs. Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines. Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual values summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
@@
In addition, ablation studies on using only value-augmented generation and on varying the number of retrieved summaries confirm ValuesRAG’s robust performance under different configurations. ValuesRAG surpasses the baselines through purely values-based generation. Meanwhile, adjusting the number of retrieved documents shows that moderate retrievals can effectively balance diversity and relevance, also maintaining high accuracy across multiple benchmarks.
@@
These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications. Our study demonstrates ValuesRAG’s robust capabilities on a global scale, also suggesting its applicability in aligning the values of diverse groups within a single country. ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes Li et al. (2024). This, in turn, facilitates the creation of fairer and more effective policies. Moreover, NGOs can leverage ValuesRAG to develop LLMs that reflect specific value orientations while maintaining strong alignment with users’ values, thereby increasing their persuasive impact. This approach benefits the promotion and spread of values that contribute to the planet’s sustainable development and the long-term well-being of human society.
2 Related Work
2.1 Evaluation of LLMs’ Cultural Bias
@@
Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue Tao et al. (2024). Unlike the more obvious safety concerns and social discrimination Liu et al. (2024b) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data. Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  Manvi et al. (2024); Durmus et al. (2024). This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  Manvi et al. (2024). Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models Gallegos et al. (2024). For example, Webster et al.  Webster et al. (2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models, while Caliskan et al. Caliskan et al. (2017) pioneered the use of word embeddings as quantitative measures of bias. More recently, Karinshak et al. Karinshak et al. (2024) introduced LLM-GLOBE, a benchmark where LLMs generate both quantitative and open-ended answers to values assessment questions, with subsequent evaluation using the LLM-as-a-Jury Protocol. These evaluation methods collectively highlight the complex nature of cultural bias in LLMs and the need for multifaceted assessment approaches.
@@
2.2 Mitigation of LLMs’ Cultural Bias
Methods like RLHF Shen et al. (2023); Ji et al. (2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not that suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al. (2024), as there are hundreds of countries and cultures in the earth. Addressing cultural biases has become a critical area of research, with various strategies being proposed to enhance cultural sensitivity in LLMs. For instance, Tao et al. Tao et al. (2024) adopted national and cultural role assignments to adjust the cultural values of LLMs, while Masoud et al. Masoud et al. (2024) developed a soft prompt tuning approach to mitigate bias. Also, Choenni and Shutova Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts. However, these approaches face significant limitations in fully capturing the complexity of cultural alignment. Tao et al.’s technique mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations. Meanwhile, Choenni and Shutova’s few-shot learning approach similarly falls short of modeling cultural alignment in all its complexity. We therefore use these methods as baselines to benchmark our proposed approach.
3 Datasets
In this section, we first introduce the World Values Survey (WVS) as our retrieval corpus (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies. Subsequently, we describe six regional test datasets (Section 3.2), which are carefully selected to ensure geographic, cultural, and demographic diversity.
3.1 Retrieval Corpus
VS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al. (2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries. WVS is selected as our retrieval corpus especially due to its numerous advantages:
@@ 3. Effective structure and large scale:
Since values evolve gradually over time, VS is conducted in waves, with each wave occurring every five years. For our study, we utilize the most recent wave, spanning from 2017 to 2022.
The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions. The value questions span 13 topics, such as social trust, post-materialism, and political interest. We randomly select 20% (52 questions) per topic for validation and use the remaining 80% (207 questions) for summary generation. The 31 demographic features, including country, sex, age, education, social class, and employment status, are used to generate demographic summaries for retrieval tasks.
Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment. Consisting four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
3.2 Test Datasets
We select six regional surveys [^2] Detailed dataset description are provided in Appendix A. to serve as test datasets based on the following criteria:
@@ 2. Temporal proximity:
The regions in our test datasets are meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring that the data accurately reflects the majority of the global population. All of them are publicly accessible and are statistically representative at national or regional levels, which guarantees their reliability and validity.
To be specific, we select the European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region. For the United States, we select the General Social Survey Davern et al. (2024), which is the most comprehensive social survey in the country. The Chinese General Social Survey Bian and Li (2012) serves as the representative dataset for China due to its comprehensive sampling methodology and scientific rigor. For India, where national survey data were largely inaccessible during our study, we use Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population. The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, is selected to represent Latin American countries, as it covers 32 countries across the region. Finally, Afrobarometer Afrobarometer (2023) is chosen as the representative dataset for Africa. A detailed summary of these datasets is presented in Table 1.
4 Methodology
@@
In this section, we present the ValuesRAG which is specifically designed to address cultural biases and enhance contextual alignment in LLM-driven scenarios through a Retrieval-Augmented Generation (RAG) approach. ValuesRAG consists of three key components: Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets (Section 4.1); Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context (Section 4.2); and Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles (Section 4.3). [^3] We provide a case study of retrieve in Appendix C. An overview of the ValuesRAG framework is provided in Figure 2.
4.1 Values and Demographic Summary Generation
@@
systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages. First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved (described in Section 3.1). In parallel, topic-based summaries and demographic summaries are generated separately. For topic-based summaries, values-related QA sets are used to produce summaries for each topic, while demographic summaries are generated using demographic-related QA sets:
@@
with $M$ denoting the total number of topics. The result, denoted as $S_{i}$, forms a comprehensive values summary for individual $i$. These generate summaries serve as structured references for retrieval in later stages and are also used to augment the validation set for evaluation.
4.2 Values Augmented Generation
@@
$C_{i}=\text{concat}\left(S_{i},\text{Q}_{i}^{\text{val},k}\right)$ (3)
where $C_{i}$ represents the combined context, $S_{i}$ is the values summary for individual $i$, and $\text{Q}_{i}^{\text{val},k}$ is the $k$-th validation question. Next, we concatenate $C_{i}$ with the demographic summary $D_{i}$ to further enhance the context, enabling the generation of responses based on both values and demographic information:
@@
Here, $A_{i}$ represents the answer generated by the function $f$, and $(C_{i},D_{i})$ embeds the augmented context $C_{i}$ and demographic information $D_{i}$ into a structured input format. Additionally, we utilize chain-of-thought prompting to enhance reasoning and emulate the behavior of the corresponding individual, ensuring responses that are contextually aligned with the values captured in the summaries and demographic characteristics.
4.3 Retrieval-based Values Alignment
To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval. The demographic data from both the train (retrieval corpus) and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model. We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
$\small\text{Sim}(E_{\text{test}},E_{\text{train}})=\frac{E_{\text{test}}\cdot E_{\text{train}}}{\|E_{\text{test}}\|\|E_{\text{train}}\|}$ (5)
Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training (retrieval corpus) demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score. The top-100 embeddings with the highest similarity scores are initially selected as candidates. Subsequently, we apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates. The reranking process evaluates the semantic relevance of the initial candidates based on a scoring function:
$\small R^{\prime}_{k}=f_{\text{rerank}}\left(E_{\text{test}},E_{C_{j}}\right),\;C_{j}\in\{C_{1},C_{2},\dots,C_{100}\},\;k\leq 100$ (6)
where $R^{\prime}_{k}$ represents the reranked summary for the $k$-th candidate selected from the final top-$k$ results, $f_{\text{rerank}}$ is the reranking function, and $E_{C_{j}}$ denotes the embedding of the $j$-th candidate initially retrieved among the 100 documents. The reranked top-$k$ summaries are then incorporated into the prompts, enriching the contextual alignment of the generated responses. In detail, for each test individual, the retrieved and reranked summaries are combined into the final prompt, and the answer is subsequently generated using the function $f_{\text{gen}}$:
@@ 4.3 Retrieval-based Values Alignment
3:for each individual $i\in\mathcal{D}$ do
@@
Model Methods EVS GSS CGSS ISD LAPOP Africa Avg.
@@
Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120

### v3

@@ Abstract
Ensuring cultural values alignment in Large Language Models (LLMs) remains a critical challenge, as these models often embed Western-centric biases from their training data, leading to misrepresentations and fairness concerns in cross-cultural applications. Existing approaches—such as role assignment and few-shot learning—struggle to address these limitations effectively due to their reliance on pre-trained knowledge, limited scalability, and inability to capture nuanced cultural values. To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with In-Context Learning (ICL) to integrate cultural and demographic knowledge dynamically during text generation. Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual. We subsequently curate several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries. We evaluate ValuesRAG using $6$ diverse regional datasets and show that it consistently outperforms baselines: including zero-shot, role-assignment, few-shot, and hybrid methods, both in main experiments and ablation settings. Notably, ValuesRAG achieves the best overall performance over prior methods, demonstrating its effectiveness in fostering culturally aligned and inclusive AI systems. Our findings underscore the potential of dynamic retrieval-based methods to bridge the gap between global LLM capabilities and localized cultural values.
@@
Introduction
The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  (Singh et al. 2024; Kharchenko et al. 2024; Hu et al. 2024). Predominantly trained on Western data sources (Achiam et al. 2023; Touvron et al. 2023; Jiang et al. 2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts. These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs (Gallegos et al. 2024; Xie et al. 2024; Potter et al. 2024; Huang, Papyshev, and Wong 2024). Despite ongoing efforts to address these issues, existing strategies often fall short. While some countries have developed localized LLMs, such as China’s ERNIE (Sun et al. 2021), ChatGLM (GLM et al. 2024), DeepSeek (Liu et al. 2024a), and South Korea’s HyperCLOVA (Yoo et al. 2024), these models also exhibit biases inherited from their respective training datasets. As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment (Gallegos et al. 2024; Xie et al. 2024; Potter et al. 2024).
@@ Introduction
Recent studies have proposed several approaches, such as role-assignment approaches (Tao et al. 2024) and few-shot learning techniques  (Choenni and Shutova 2024), to mitigate these cultural biases. However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
@@
To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL) to dynamically incorporate cultural knowledge during text generation (see Figure 1). Our framework leverages the World Values Survey (WVS) dataset (Haerpfer et al. 2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies. First, we specifically generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel. After constructing the knowledge base, we retrieve the top-$100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected. Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies reasoning grounded in the retrieved values to generate the final answer to the question.
@@
We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method (Tao et al. 2024), (3) few-shot learning (Choenni and Shutova 2024), and (4) a hybrid method combining both (1) and (2). To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs. Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines. Compared to previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual values summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
@@
In addition, ablation studies on varying the number of retrieved summaries and on using only value-augmented generation confirm ValuesRAG’s robust performance under different configurations. Adjusting the number of retrieved documents shows that moderate retrievals can balance diversity and relevance, also maintaining high accuracy across multiple benchmarks. Meanwhile, ValuesRAG surpasses the baselines through purely values-based generation.
@@
These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications. Our study demonstrates ValuesRAG’s robust capabilities on a global scale, also suggesting its applicability in aligning the values of diverse groups within a single country. ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes (Li, Gong, and Jiang 2024). This, in turn, facilitates the creation of fairer and more effective policies. Moreover, NGOs can leverage ValuesRAG to develop LLMs that reflect specific value orientations while maintaining strong alignment with users’ values, thereby increasing their persuasive impact. This approach benefits the promotion and spread of values that contribute to the planet’s sustainable development and the long-term well-being of human society.
Related Work
Evaluation of LLMs’ Cultural Bias
@@
Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue (Tao et al. 2024). While safety concerns and social discrimination in language models are typically explicit and well-recognized (Liu et al. 2024b), cultural biases often manifest subtly, reflecting dominant cultural perspectives embedded within training data. Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  (Manvi et al. 2024; Durmus et al. 2024). This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  (Manvi et al. 2024). Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models (Gallegos et al. 2024). For instance, Caliskan et al. (2017) pioneered the use of word embeddings as quantitative measures of bias, while Webster et al.  (2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models. More recently, Karinshak et al. (2024) introduced LLM-GLOBE, a benchmark where LLMs generate both quantitative and open-ended answers to values assessment questions, with subsequent evaluation using the LLM-as-a-Jury Protocol. These evaluation methods collectively highlight the complex nature of cultural bias in LLMs and the need for multifaceted assessment approaches.
@@
Mitigation of LLMs’ Cultural Bias
Techniques such as Reinforcement Learning from Human Feedback (RLHF) (Shen et al. 2023; Ji et al. 2024) are commonly employed for aligning LLMs with human values. However, these single-dimensional alignment methods are insufficient for mitigating cultural bias, as cultural values are inherently diverse, dynamic, and context-dependent, varying significantly across different regions and societies (Huang, Papyshev, and Wong 2024). Addressing cultural biases has become a critical area of research, with various strategies being proposed to enhance cultural sensitivity in LLMs. For instance, Tao et al.  (2024) adopted national and cultural role assignments to adjust the cultural values of LLMs, while Masoud et al. (2024) developed a soft prompt tuning approach to mitigate bias. Moreover, Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts. However, these approaches face significant limitations in fully capturing the complexity of cultural alignment. Tao et al.’s approach (Tao et al. 2024) mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations. Meanwhile, Choenni and Shutova’s few-shot learning approach (Choenni and Shutova 2024) similarly falls short of modeling cultural alignment in all its complexity. We therefore use these methods as baselines to benchmark our proposed approach.
Datasets
In this section, we first introduce the World Values Survey (WVS) as our retrieval corpus, highlighting its extensive coverage, global representativeness, and relevance for values-related studies. Subsequently, we describe six regional test datasets, which are carefully selected to ensure geographic, cultural, and demographic diversity.
Retrieval Corpus
WVS [^1] https://www.worldvaluessurvey.org/wvs.jsp (Haerpfer et al. 2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries. WVS is selected as our retrieval corpus especially due to its numerous advantages:
@@ 3. Effective structure and large scale:
Since values evolve gradually over time, WVS is conducted in waves, with each wave occurring every five years. For our study, we utilize the most recent wave, spanning from 2017 to 2022.
The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions. The value questions span 13 topics, such as social trust, post-materialism, and political interest, as shown in Table 2. It covers most dimensions of values, allowing for a comprehensive and accurate measure of each respondent’s values. We randomly select 20% (52 questions) per topic for validation and use the remaining 80% (207 questions) for summary generation. The 31 demographic features, including country, sex, age, education, social class, and employment status, are used to generate demographic summaries for retrieval tasks.
Figure 2: Distribution of countries in the WVS dataset. The WVS dataset covers surveys conducted in over $120$ countries across all major regions, providing broad geographic and demographic representation for building a reliable retrieval corpus in RAG-based frameworks.
Topic Count
Social Values, Norms, Stereotypes 45
Happiness and Wellbeing 11
Social Capital, Trust and Organizational Membership 47
Economic Values 6
Perceptions of Corruption 9
Perceptions of Migration 10
Perceptions of Security 21
Index of Postmaterialism 6
Perceptions about Science and Technology 6
Religious Values 12
Ethical Values 23
Political Interest and Political Participation 35
Political Culture and Political Regimes 25
Test Datasets
We select six regional surveys to serve as test datasets based on the following criteria:
@@ 2. Temporal proximity:
The regions in our test datasets are meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring that the data accurately reflects the majority of the global population. All of them are publicly accessible and are statistically representative at national or regional levels, which guarantees their reliability and validity. Also, these test sets include both values-related questions and demographic characteristics. The demographic characteristics are used to generate summaries, serving as retrieval targets for RAG. The values-related questions are utilized as test questions to calculate accuracy (ACC). [^2] A detailed description of the method for computing accuracy (ACC) is provided in Section Experiments - Evaluation Metrics. The specific datasets used in our evaluation, along with their characteristics, are described below:
EVS
GSS
CGSS
ISD
To ensure that our experiment covers as much of the world’s population as possible, we made efforts to include India within the scope of our test set. However, we were unable to obtain data from several government surveys in India, thus we used data published by the Pew Research Center instead. The Pew Research Center’s India Survey Dataset [^6] https://www.pewresearch.org/dataset/india-survey-dataset is a comprehensive resource that captures the perspectives of 29,999 Indian adults on various aspects of society, including religious beliefs and practices, identity, nationalism, and societal tolerance. Conducted through face-to-face interviews between November 17, 2019, and March 23, 2020, the survey encompassed participants from diverse religious backgrounds, such as Hindus, Muslims, Sikhs, Christians, Buddhists, Jains, and others. This dataset covers 33 values-related questions and 23 demographic characteristics.
LAPOP
Africa
Figure 3: Overview of the proposed ValuesRAG framework for cultural alignment. The framework comprises four key stages: (1) Dataset preprocessing to separate training and test QA pairs from WVS and regional datasets, (2) Topic-wise summary generation using LLMs for each individual, (3) Aggregation of topic summaries into comprehensive individual values profiles, and (4) Retrieval-Augmented Generation that retrieves and reranks relevant value summaries based on demographic similarity to guide final response generation.
Methodology
@@
In this section, we present the ValuesRAG which is specifically designed to address cultural biases and enhance contextual alignment in LLM-driven scenarios through a Retrieval-Augmented Generation (RAG) approach. ValuesRAG consists of three key components: (1) Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets; (2) Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context; and (3) Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles. An overview of the ValuesRAG framework is provided in Figure 3.
Values and Demographic Summary Generation
@@
To systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages. First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved, as described in previous section. In parallel, topic-based summaries and demographic summaries are generated separately. For topic-based summaries, values-related QA sets are used to produce summaries for each topic, while demographic summaries are generated using demographic-related QA sets:
@@
with $M$ denoting the total number of topics. The result, denoted as $S_{i}$, forms a comprehensive values summary for individual $i$. These generate summaries serve as structured references for retrieval in later stages and are used to augment the validation set for evaluation as well.
Values Augmented Generation
@@
$C_{i}=\text{concatenate}\left(S_{i},\text{Q}_{i}^{\text{val},k}\right)$ (3)
where $C_{i}$ represents the combined context, $S_{i}$ is the values summary for individual $i$, and $\text{Q}_{i}^{\text{val},k}$ is the $k$-th validation question. We subsequently concatenate $C_{i}$ with the demographic summary $D_{i}$ to further enhance the context, enabling the generation of responses based on both values and demographic information:
@@
Here, $A_{i}$ represents the answer generated by the function $f$, and $(C_{i},D_{i})$ embeds the augmented context $C_{i}$ and demographic information $D_{i}$ into a structured input format. Additionally, we utilize Chain-of-Thought (CoT) prompting (Wei et al. 2022) to enhance reasoning and emulate the behavior of the corresponding individual, ensuring responses that are contextually aligned with the values captured in the summaries and demographic characteristics.
Retrieval-based Values Alignment
To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval. The demographic data from both the train (retrieval corpus) and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a sentence-transformer-based model (Reimers and Gurevych 2019). [^9] A detailed description of the model used is provided in  Section Experiments - Models Used. We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
$\text{Sim}(E_{\text{test}},E_{\text{train}})=\frac{E_{\text{test}}\cdot E_{\text{train}}}{\|E_{\text{test}}\|\|E_{\text{train}}\|}$ (5)
$E_{\text{test}}$ and $E_{\text{train}}$ specifically represent the embeddings of the test and training (retrieval corpus) demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score. The top-100 embeddings with the highest similarity scores are initially selected as candidates. We subsequently apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates.
The reranking process evaluates the semantic relevance by passing each candidate summary embedding $E_{C_{j}}$ along with the test individual’s embedding $E_{\text{test}}$ through a neural reranker $f_{\text{rerank}}$, which outputs a relevance score $s_{j}$:
$s_{j}=f_{\text{rerank}}\left(E_{\text{test}},E_{C_{j}}\right),\quad j\in\{1,2,\dots,100\}$ (6)
We then sort the candidate summaries based on these scores in descending order and select the top-$k$ summaries with the highest scores as the final reranked set $R^{\prime}_{k}$:
$R^{\prime}_{k}=\text{Top-}k\left(\{C_{j}\}_{j=1}^{100},\{s_{j}\}_{j=1}^{100}\right)$ (7)
The reranked top-$k$ summaries $R^{\prime}_{k}$ are incorporated into the prompts, enriching the contextual alignment of the generated responses. In detail, for each test individual, the retrieved and reranked summaries are combined into the final prompt, and the answer is subsequently generated using the function $f_{\text{gen}}$:
@@
3:for each individual $i$ in $\mathcal{D}$ do
@@
Experiments
Setup
Models Used.
We utilize GPT-4o-mini (Achiam et al. 2023) and Gemini-1.5-Flash (Team et al. 2024) for our generation tasks, which are accessed via APIs. We set the temperature parameter of these models to $0.7$ to achieve a balance between coherence and creativity. For the retrieval task, we employed the E5 (base) model (Wang et al. 2022), which generates embeddings and retrieves the top 100 most relevant summaries of values based on cosine similarity.
To refine this retrieval, we apply a reranker using the GTE-multilingual-reranker-base model (Zhang et al. 2024). Each retrieved candidate is paired with the test individual’s demographic summary and passed as a sentence pair to the reranker. The model outputs a relevance score for each pair, which we use to rank all candidates. The top-k summaries with the highest scores are then selected as the final set used for generation. This two-stage process: retrieval followed by reranking, ensures both semantic similarity and contextual relevance.
Prompts Used
We provide the prompts designed for various components of the ValuesRAG, including prompts for performing question answering tasks, as well as for generating values and demographic summaries below:
Baseline Methods and Implementation.
Our baseline methods include: (1) Zero-shot inference, (2) the role-assignment-only approach (Tao et al. 2024), (3) a few-shot learning method (Choenni and Shutova 2024), and (4) a hybrid method that combines both (1) and (2). For the role-assignment baseline, we specifically use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data. For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts. The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set. Additionally, We use ValuesRAG with $k=3$ retrieved summaries—chosen to provide a good balance between retrieval diversity and contextual relevance. [^10] A detailed analysis of varying $k$ and its implications on model performance is provided in Ablation Study - Varying the Number of Retrieved Summaries.
Evaluation Metrics.
We utilize accuracy as the primary evaluation metric, following previous work (Choenni and Shutova 2024), by converting multiple-choice responses into a binary format for consistency and simplicity. Specifically, we transform each original response $r_{i}$, ranging on a scale (e.g., Likert scale from 1 to 10), into a binary value $b_{i}$ indicating disagreement or agreement as follows. Here, $m$ denotes the midpoint of the response scale:
$b_{i}=\begin{cases}0,&\text{if }r_{i}\leq m\quad\text{(disagree)}\\[6.0pt] 1,&\text{if }r_{i}>m\quad\text{(agree)}\end{cases}$ (9)
This categorization metrics effectively captures distinct answer patterns and aligns naturally with values-related questions often posed through Likert-scale formats.
Model Methods EVS GSS CGSS ISD LAPOP Africa Avg. Accuracy
@@
Role-Assignment (2024) 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
Few-Shot Learning (2024) 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120